OpenStudy (anonymous):

Slope help needed- a medal is rewarded Launch Area: (1, 2) Point A: (-6, 5) Point B: (-9, 8) Point C: (7, -3)

OpenStudy (anonymous):

Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B. Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C.

OpenStudy (anonymous):

The slope for Launch to A is -3/7 The slope for A to B is 13/3 The slope for B to C is 5/16 I started with these on my own but they didn't look right.

OpenStudy (anonymous):

@jim_thompson5910 @UnkleRhaukus @.Sam. @ash2326 @jagatuba @Preetha

OpenStudy (anonymous):

@whalexnuker You did it right for Launch to A but err in calculating slope for A to B and for B to C

OpenStudy (anonymous):

for A to B the slope will be -1 and for B to C it will be -11/16

OpenStudy (anonymous):

slope between two points,lets say (x1,y1) and (x2,y2) in 2-D is given by (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Huh? "but err"? y2-y1/x2-x1 = 8-5/-9-(-6) = 3/-3 = -1 A: -6, 5 B: -9, 8 Are you sure it's -1? Because in the beginning, I added 6 since two negatives make a positive

OpenStudy (anonymous):

y2-y1/x2-x1 = -3-8/7-(-9) = -11/16 Point B: (-9, 8) Point C: (7, -3) Still, are you sure? Because I added (again) because two positives make a negative

OpenStudy (anonymous):

Hey adding two negative integers will always give negative integer. Multiplying two negative values makes the product positive

OpenStudy (anonymous):

Yup I am damn sure about it.. Don't u get it?

OpenStudy (anonymous):

@whalexnuker ??

OpenStudy (anonymous):

uh, no. Adding two negatives give you a positive e_e @quantun